A 7.09 µF capacitor and a 7.39 mH inductor are connected in series with an AC power source that has a frequency of 3.45 x103 Hz and a peak voltage of 65 V. Take the initial time t as zero when the instantaneous voltage equals zero. Determine the instantaneous current when t = 2.13x 10-4 s.
Answer #1

Net reactance X = [2pifL - 1/(2pifC]

= 2pi*3450*7.39e-3 -1/(2pi*3450*7.09e-6)

= 153.7 ohm

Here voltage V = 65*sin (2pi*3450*t)

current I = (65/X)*sin (2pi*3450*t -pi/2)

current I = 65*sin (2pi*3450*2.13e-4 - pi/2)/153.7

= 0.0402 A answer